OpenStudy (yuffanhza):

Consider the vector-valued function r(f) = e^-t i+ e^-t cos t j + e^-t sin t k. a) Find r'(t) and |r'(t)| b) Evaluate integrate x^2 ds, where C is the graph of r(t), 0 < t

OpenStudy (eliesaab):

To find r'(t), find the derivative of each componnent

OpenStudy (eliesaab):

\[ r'(t)=\left\{-e^{-t},-e^{-t} \sin (t)-e^{-t} \cos (t),e^{-t} \cos (t)-e^{-t} \sin (t)\right\} \]

OpenStudy (eliesaab):

\(|r'(t)|\) is the length of the vector \( r'(t)\)

OpenStudy (eliesaab):

You find \[ |r'(t)|=\sqrt{e^{-2 t}+\left(-e^{-t} \sin (t)-e^{-t} \cos (t)\right)^2+\left(e^{-t} \cos (t)-e^{-t} \sin (t)\right)^2} \]

OpenStudy (eliesaab):

Expand the above and simplify

OpenStudy (eliesaab):

You get \[ |r'(t)|=\sqrt{3} \sqrt{e^{-2 t}}=\sqrt{3} e^{-t} \] You hsve to fill in the details

OpenStudy (eliesaab):

You need now to compute \[ \int_0^{\ln 3} x^2 ds=\int_0^{\ln 3} x(t)^2 |r'(t)| dt \] Fill in the detaila

OpenStudy (eliesaab):

details

OpenStudy (eliesaab):

\[ x(t)=e^{-t} \]

OpenStudy (eliesaab):

The rest is easy